OpenStudy (anonymous):

Can somebody help me with this!(will give medals) 2/3 x minus 5/8 x =26

OpenStudy (anonymous):

I need help! ive been stuck for more then two hours!

OpenStudy (richyw):

\[\frac{2}{3}x-\frac{5}{8} x = 26\]

OpenStudy (anonymous):

Thats right.

OpenStudy (richyw):

do you know how to add and subtract fractions?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

cdm is 24

OpenStudy (richyw):

ok so what you can do is factor out the x: \[\left(\frac{2}{3}-\frac{5}{8}\right)x=26\]

OpenStudy (anonymous):

Ok

OpenStudy (richyw):

yup, 24 is the cdm so

OpenStudy (anonymous):

so it will be -3/24 x =26

OpenStudy (anonymous):

Ohh what did i do wrong?

OpenStudy (richyw):

\[\frac{16-15}{24}=\frac{1}{24}\]

OpenStudy (anonymous):

one second while i get a pen

OpenStudy (anonymous):

ok i'm writng this down

OpenStudy (richyw):

ok so you have \[\frac{1}{24}x=26\]

OpenStudy (anonymous):

ok one second i am writing all of this in order

OpenStudy (richyw):

now you want to isolate x, so divide both sides by 1/24. Of course this is equivalent to multiplying both sides by 24.\[x=26\times24\]\[x=624\]

OpenStudy (anonymous):

oh ok

OpenStudy (richyw):

you can use wolfram alpha to check your answers: http://www.wolframalpha.com/input/?i=2%2F3+x+minus+5%2F8+x+%3D26 but of course you want to do them yourself, because there is no wolfram alpha on the exam!

OpenStudy (anonymous):

How would you check if the answer was right!

OpenStudy (richyw):

click that link, and scroll down to where it says "solution"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That's realy helpful! Thank you!

OpenStudy (anonymous):

I will keep on studing hard!

OpenStudy (richyw):

right on!

OpenStudy (anonymous):

:)